OpenStudy (anonymous):

Determine the fraction that is equivalent to the limit of the sequence: 0.6,0.66,0.666,0.6666

OpenStudy (anonymous):

do u need the sum of all these numbers...?

OpenStudy (anonymous):

let S= 0.6+0.66+0.666+0.6666..... S=6/9*(0.9+0.99 +0.999 +0.9999 ...) S=6/9*((1-0.1)+(1-0.01) +0.(1-0.001) +1-0.0001 ...)

OpenStudy (anonymous):

can u do further

OpenStudy (anonymous):

or is that just u need 2/3=.666666666666666.......

OpenStudy (anonymous):

yeah the fraction is 2/3, but how do you get that?